OpenStudy (anonymous):

GIVING MEDAL! if f(x)=2x+1and g(x)= x^2 -7, find (f ' g)(x)

OpenStudy (anonymous):

f(g(x))=f(x^2-7)=2*(x^2-7)+1

OpenStudy (anonymous):

A. 2x^3+2x-14x-7 B. 2x^2-13 C. 2x^3-2x-7x-7 D. 3x^2+7x-6

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Sometimes I feel sorry for humanity.